Prove that for every positive integer n. 1.2.3+2.3.4++ n(n+1)(n+2) = n(n+1)(n+2)(n+3)/4.
1
Expert's answer
2021-08-12T04:47:02-0400

Let "P(n)" be the proposition that for every positive integer "n,"


"1\cdot2\cdot3+2\cdot3\cdot4+...+n(n+1)(n+2)=\dfrac{n(n+1)(n+2)(n+3)}{4}"



BASIC STEP: "P(1)" is true, because "1\cdot2\cdot3=\dfrac{1(1+1)(1+2)(1+3)}{4}"

INDUCTIVE STEP: For the inductive hypothesis we assume that "P(k)" holds for an arbitrary

positive integer "k." That is, we assume that


"1\cdot2\cdot3+2\cdot3\cdot4+...+k(k+1)(k+2)=\dfrac{k(k+1)(k+2)(k+3)}{4}"

Under this assumption, it must be shown that "P(k+1)" is true, namely, that


"1\cdot2\cdot3+2\cdot3\cdot4+...+k(k+1)(k+2)+(k+1)(k+2)(k+3)="

"=\dfrac{(k+1)(k+2)(k+3)(k+4)}{4}"

is also true.


"1\cdot2\cdot3+2\cdot3\cdot4+...+k(k+1)(k+2)+(k+1)(k+2)(k+3)"

"=\dfrac{k(k+1)(k+2)(k+3)}{4}+(k+1)(k+2)(k+3)"

"=\dfrac{(k+1)(k+2)(k+3)(k+4)}{4}"

This equation shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we

know that "P(n)" is true for all positive integers "n." That is, we have proven that


"1\cdot2\cdot3+2\cdot3\cdot4+...+n(n+1)(n+2)=\dfrac{n(n+1)(n+2)(n+3)}{4}"

for all positive integers "n."